Solve the system using substitution.
4x+y=2
3y+2x=-1
Thanks for helping me out :)

Answer by ankor@dixie-net.com(22740)

Solve the system using substitution.
4x + y = 2
3y + 2x =-1
:
We are going to use the 1st equation for substitution in the 2nd equation
put it in the slope/intercept form
4x + y = 2
Subtract 4x from each side
y = -4x + 2
:
In the 2nd equation, replace y with (-4x+2)
3(-4x+2) + 2x = -1
-12x + 6 + 2x = -1
-12x + 2x = -1 - 6
-10x = -7
x = -7/-10
x = +.7
:
y = -4x + 2
replace x with .7
y = -4(.7) + 2
y = -2.8 + 2
y = -.8
;
:
Check solutions in the 2nd equation
3y + 2x = -1
replace x and y
3(-.8) + 2(.7) = -1
-2.4 + 1.4 = - 1
-1 = -1; confirms our solution of x=.7; y=-.8